I have another here I do not understand how to do. Would someone show me how / Thanks for any help at all.
Find the reciprocal of 1/3 (-6-10).Write your answer in simplest form.

Answer by checkley71(8403)

(1/3)(-6-10)=1/3(-16)=-16/3 thus the reciprical=-3/16